OpenStudy (anonymous):

Parallelogram FGHI on the coordinate plane below represents the drawing of a horse trail through a local park: In order to build a scale model of the trail, the drawing is enlarged on the coordinate plane. If two corners of the trail are at point A (0, 4) and point D (-6, -5), what is another point that could represent a corner of the trail?

OpenStudy (anonymous):

http://gyazo.com/8b011e4b8f8a541fb7d15e77b724b842

OpenStudy (perl):

are A and D opposite corners ?

OpenStudy (anonymous):

A. (-2, -6) B. (6, -5) C. (11, 4) D. (-17, 5)

OpenStudy (anonymous):

I don't know, tbh. I need help with this. @perl

OpenStudy (bibby):

since they're parallelograms and A corresponds with F and D corresponds with I, I think we can use narrow our choices down to b or c because they have the same y coordinates as A and B

OpenStudy (bibby):

also, the new parallelogram is scaled by a factor of 3. the lengths of the original are 4. 4*3 = 12. so point B would be at (0+12,4) = (12,4) and point C would be at (-6+12,-5) = (6,-5)